# TestNUC: Enhancing Test-Time Computing Approaches through Neighboring Unlabeled Data Consistency

### v1

@@ Abstract
Test-time computing approaches, which leverage additional computational resources during inference, have been proven effective in enhancing large language model performance. This work introduces a novel, linearly scaling approach, TestNUC, that improves test-time predictions by leveraging the local consistency of neighboring unlabeled data—it classifies an input instance by considering not only the model’s prediction on that instance but also on neighboring unlabeled instances. We evaluate TestNUC across eight diverse datasets, spanning intent classification, topic mining, domain discovery, and emotion detection, demonstrating its consistent superiority over baseline methods such as standard prompting and self-consistency. Furthermore, TestNUC can be seamlessly integrated with existing test-time computing approaches, substantially boosting their performance. Our analysis reveals that TestNUC scales effectively with increasing amounts of unlabeled data and performs robustly across different embedding models, making it practical for real-world applications. Our code is available at https://github.com/HenryPengZou/TestNUC.
@@
TestNUC: Enhancing Test-Time Computing Approaches through Neighboring Unlabeled Data Consistency
@@ 1 Introduction
Test-time computing approaches, which leverage additional computational resources during inference to enhance performance, have gained increasing attention in the era of large language models (LLMs) Snell et al. (2024); Dong et al. (2024). There are two primary strategies for modifying an LLM’s distribution at test time: (1) at the input level: augmenting the prompt with additional tokens (e.g., few-shot in-context learning Mosbach et al. (2023)); or (2) at the output level: sampling multiple candidate answers and aggregating them (e.g., self-consistency Wang et al. (2023b), best-of-N Beeching et al. (2024)). Despite demonstrating promising capabilities, input-augmentation approaches incur a computational cost that scales quadratically with the number of added tokens in the prompt, making them more computationally expensive than output-sampling methods. Meanwhile, output-sampling approaches typically overlook the potential of large amounts of unlabeled data that are often available in real-world settings Berthelot et al. (2019); Sohn et al. (2020); Zou and Caragea (2023); Zou et al. (2025).
@@
Motivated by these findings, we propose TestNUC, a simple yet effective approach that enhances test-time LLM predictions by leveraging neighboring unlabeled data consistency. Concretely, TestNUC consists of two key steps: ❶ Neighbor Retrieval, where we identify the top-K nearest unlabeled neighbors of a test sample based on feature similarity; and ❷ Collaborative Prediction, where the LLM generates predictions for both the test sample and its retrieved neighbors, which are then aggregated to obtain the final answer. The intuition behind TestNUC is that samples in close proximity within the embedding space are likely to share similar labels. By incorporating predictions of nearby unlabeled samples, the LLM can exploit the consistency of local data structures to better contextualize and refine its decision-making, effectively using unlabeled examples as an auxiliary signal to boost test-time performance while reducing noise and uncertainty Van Gansbeke et al. (2020); Zhou et al. (2024).
@@ 3.1 Framework Overview
Algorithm 1 TestNUC algorithm.
2:  $\mathcal{N}=\arg\!\operatorname{top}_{K}\{\mathcal{S}_{f}(x_{0},x_{i})|\;i=0,\dots,N\}$Mine top-$K$ neighbors based on similarity, note that test sample $x_{0}$ is included}
3:  for $k=1$ to $K$ do
[…]
5:     $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}w_{k}=\operatorname{sim}(z_{0},z_{\mathcal{N}_{k}})$Compute neighbor weights based on similarity}
[…]
7:  end for
[…]
9:  Return $y_{\mathrm{final}}$
@@ Appendix B Prompt Template
Prompt Template of FFA Instruction: Please select a label from the provided options for the following testing samples and also show your confidence in the label assignment by providing a probability between 0 and 1. Label Options: [A List of Labels]. == Testing Samples == [Testing Samples]

### v2

# TestNUC: Enhancing Test-Time Computing Approaches and Scaling through Neighboring Unlabeled Data Consistency

@@ Abstract
Test-time computing approaches, which leverage additional computational resources during inference, have been proven effective in enhancing large language model performance. This work introduces a novel, linearly scaling approach, TestNUC, that improves test-time predictions by leveraging the local consistency of neighboring unlabeled data-it classifies an input instance by considering not only the model’s prediction on that instance but also on neighboring unlabeled instances. We evaluate TestNUC across eight diverse datasets, spanning intent classification, topic mining, domain discovery, and emotion detection, demonstrating its consistent superiority over baseline methods such as standard prompting and self-consistency. Furthermore, TestNUC can be seamlessly integrated with existing test-time computing approaches, substantially boosting their performance. Our analysis reveals that TestNUC scales effectively with increasing amounts of unlabeled data and performs robustly across different embedding models, making it practical for real-world applications. Our code is available at https://github.com/HenryPengZou/TestNUC.
@@
TestNUC: Enhancing Test-Time Computing Approaches and Scaling through Neighboring Unlabeled Data Consistency
@@ 1 Introduction
Test-time computing approaches, which leverage additional computational resources during inference to enhance performance, have gained increasing attention in the era of large language models (LLMs) Snell et al. (2024); Dong et al. (2024). There are two primary strategies for modifying an LLM’s distribution at test time: (1) at the input level: augmenting the prompt with additional tokens (e.g., few-shot in-context learning Mosbach et al. (2023)); or (2) at the output level: sampling multiple candidate answers and aggregating them (e.g., self-consistency Wang et al. (2023b), best-of-N Beeching et al. (2024)). Despite demonstrating promising capabilities, input-augmentation approaches incur a computational cost that scales quadratically with the number of added tokens in the prompt, making them more computationally expensive than output-sampling methods. Meanwhile, output-sampling approaches typically overlook the potential of large amounts of unlabeled data that are often available in real-world settings Berthelot et al. (2019); Sohn et al. (2020); Zou and Caragea (2023); Zou et al. (2025); Gu et al. (2025).
@@
Motivated by these findings, we propose TestNUC, a simple yet effective approach that enhances test-time LLM predictions by leveraging neighboring unlabeled data consistency. Concretely, TestNUC consists of two key steps: ❶ Neighbor Retrieval, where we identify the top-K nearest unlabeled neighbors of a test sample based on feature similarity; and ❷ Collaborative Prediction, where the LLM generates predictions for both the test sample and its retrieved neighbors, which are then aggregated to obtain the final answer. The intuition behind TestNUC is that samples in close proximity within the embedding space are likely to share similar labels. By incorporating predictions of nearby unlabeled samples, the LLM can exploit the consistency of local data structures to better contextualize and refine its decision-making, effectively using unlabeled examples as an auxiliary signal to boost test-time performance while reducing noise and uncertainty Van Gansbeke et al. (2020); Zhou et al. (2024); Wang et al. (2025).
@@ 3 Method
Algorithm 1 TestNUC algorithm.
2:  $z_{0}=f(x_{0})$, $\mathcal{Z}=\{z_{i}=f(u_{i})\}_{i=1}^{N}$Extract embeddings for test sample and unlabeled data}
3:  $\mathcal{N}=\arg\!\operatorname{top}_{K}\{\mathcal{S}_{f}(x_{0},x_{i})|\;i=0,\dots,N\}$Mine top-$K$ neighbors based on similarity, note that test sample $x_{0}$ is included}
4:  for $k=1$ to $K$ do
[…]
6:     $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}w_{k}=\operatorname{sim}(z_{0},z_{\mathcal{N}_{k}})$Compute neighbor weights based on similarity}
[…]
8:  end for
[…]
10:  Return $y_{\mathrm{final}}$
@@ 8 Limitation
Acknowledgements
This work is supported in part by NSF under grants III-2106758, and POSE-2346158.
@@ Appendix B Prompt Template
Prompt Template Instruction: Please select a label from the provided options for the following testing samples and also show your confidence in the label assignment by providing a probability between 0 and 1. Label Options: [A List of Labels]. == Testing Samples == [Testing Samples]
@@ Appendix C LLM Predictions Can Be Inaccurate and Unstable
Appendix D Robustness in Adversarial Scenarios
In real-world applications, suitable in-distribution data may not always be available, and retrieved samples could be out-of-distribution. This section demonstrates the robustness of our TestNUC method in such adversarial scenarios. We conducted an adversarial experiment by replacing in-distribution samples in the Banking dataset with out-of-distribution samples (i.e., outliers) from the Reddit dataset. To create sufficiently challenging scenarios, we replaced 60% and 75% of the in-distribution samples with OOD samples. As shown in the Table 6, even with 60%–75% OOD samples present, TestNUC still significantly improves baseline performance across many cases in these highly noisy scenarios, demonstrating its robustness and effectiveness.
Furthermore, when using Weighted Majority Voting (WMV) instead of Naive Majority Voting (NMV), TestNUC’s performance and robustness can be further enhanced. This is because OOD samples are likely to have lower semantic similarity with the test sample compared to in-distribution samples, and thus the model can assign lower weights to OOD samples when using WMV. This is also consistent with our previous findings that WMV outperforms NMV in most cases.
OOD Ratios 60% 75%
# Neighbors NMV WMV NMV WMV
0 63.7±0.6 63.7±0.6 63.7±0.6 63.7±0.6
3 67.9±0.7 68.2±0.9 68.3±0.7 69.2±0.9
5 69.6±0.8 70.6±0.8 70.3±0.7 71.1±0.6
10 72.3±0.8 72.8±0.6 73.7±0.7 74.0±0.5
20 74.7±0.3 75.0±0.4 73.3±0.7 73.3±0.8
30 75.9±0.7 75.6±0.6 72.1±0.6 72.8±0.7
40 75.7±0.8 76.3±0.6 69.1±0.8 70.8±0.9
50 74.7±0.6 75.2±0.6 67.0±0.7 69.6±0.8
60 73.5±0.6 73.7±0.5 62.0±0.8 66.0±0.7
80 71.3±0.6 72.1±0.7 54.9±0.6 61.7±0.8
100 67.3±0.7 70.1±0.7 48.8±0.6 59.0±0.7
Table 6: Performance of TestNUC under adversarial conditions with out-of-distribution (OOD) samples.
Appendix E Runtime and Cost Analysis
This section presents an estimated analysis of the trade-off between gains in accuracy (average on all 8 tasks) and the compute/runtime cost required for the approach in different settings. The following table summarizes the estimated runtime and cost per sample using GPT-4o-mini. The number of retrieved neighbors and sampled candidate answers is set to 10 by default, unless otherwise specified.
We include two variants of TestNUC: TestNUC and TestNUC-S. TestNUC-S is an efficient implementation of TestNUC that pre-computes and stores the embeddings and predictions of previously queried or seen samples. Thus, when a new sample arrives, TestNUC-S can simply retrieve the embeddings and predictions of the nearest neighbors from the stored set and use them to generate the final label without additional LLM calls—where the runtime cost for retrieval is negligible compared to querying the LLM. As shown in Table 7, when increasing the number of retrieved neighbors (K=5 to K=50), TestNUC can greatly improves performance although at the cost of increased runtime and cost. When using the same number of retrieved neighbors (K=10), TestNUC is more efficient than the best-performing baseline KNN-ICL-P, which incurs a computational cost that scales quadratically with K, whereas TestNUC incurs only a linear cost. Moreover, the efficient implementation TestNUC-S can significantly reduce runtime cost and is also a very practical solution for real-world applications, as storing the embeddings and queries of previously queried samples is both quite common and cheap in practice.
Runtime (s) Cost (\$) Performance
Standard Prompting 0.6820 \$0.000028 0.613
w. TestNUC (K=5) 3.4100 \$0.000140 0.648
w. TestNUC (K=10) 6.8200 \$0.000280 0.660
w. TestNUC (K=50) 34.1000 \$0.001400 0.676
w. TestNUC-S (K=5) 0.6842 \$0.000028 0.648
w. TestNUC-S (K=10) 0.6843 \$0.000028 0.660
w. TestNUC-S (K=50) 0.6847 \$0.000028 0.676
KNN-ICL 0.6972 \$0.000085 0.630
KNN-ICL-P 7.5510 \$0.000371 0.657
Self-Consistency 6.8200 \$0.000280 0.625
Best-of-N 6.8200 \$0.000280 0.627
Table 7: Runtime and cost analysis. TestNUC improves performance with more neighbors at higher cost, while TestNUC-S achieves the same gains with negligible runtime by reusing stored embeddings and predictions.